OpenStudy (anonymous):

dentify the 15th term of the geometric sequence 52, 39, 29.25 ... Select one: A. a15 ≈ 0.19 B. a15 ≈ 0.13 C. a15 ≈ 0.93 D. a15 ≈ 0.69

OpenStudy (anonymous):

to find the nth term in the sequence you will have to use the formula an = a1r^(n-1) In your case r = 39/52 and a1 = 52. so a15 = 52 * (39/52)^14

OpenStudy (anonymous):

Is it B?